OpenStudy (anonymous):

2(ax-by)+a+4b=0 2(bx+ay)+b-4a=0 find value of x & y by cross multiplication

OpenStudy (anonymous):

pl solve my problem

OpenStudy (anonymous):

i get \[x=-\frac{1}{2}\]

OpenStudy (anonymous):

\[(2ax-by)=-a-4b\] \[a(bx+ay)=4a-b\]

OpenStudy (anonymous):

multiply the first equation by b, the second one by a get \[2(a^2x-aby)=-a^2-4ab\] \[2(b^x+aby)=4ab-b^2\]

OpenStudy (anonymous):

ok lots of typos here but the idea is correct second line above should be \[2(b^2x +aby)=4ab-b^2\]

OpenStudy (anonymous):

now add first and second equation to get \[2(a^2+b^2)x=-a^2-b^2\] \[2(a^2+b^2)x=-(a^2+b^2)\] \[x=-\frac{a^2+b^2}{2(a^2+b^2)}=-\frac{1}{2}\]

OpenStudy (anonymous):

assuming of course that a and b are not both zero.

OpenStudy (anonymous):

and assuming that you come back and check this question

OpenStudy (a_clan):

x -y 1 ------------------- = ----------------- = --------------- -2b(b-4a) - 2a(a+4b) 2a(b-4a)-2b(a+4b) 2a.2a - 2b.-2b

OpenStudy (anonymous):

Solve both sides for x & then set equal to each other to end with y = 2. Substitute 2 in for y in either equation and you should get x = -1/2